OpenStudy (stamp):

Calculus III: Lines, planes, and surfaces in space. Find the angel and line of intersection between two planes.

OpenStudy (stamp):

Find the angle and the line of intersection between\[5x-3y+z=0\]\[x+4y+7z=1\] @B25

OpenStudy (stamp):

\[\frac{x-\frac{3}{23}}{-25}=\frac{y-\frac{5}{23}}{-34}=\frac{z}{23}\]

OpenStudy (anonymous):

I can help you with that

OpenStudy (anonymous):

\[x=3/23-25t, y=5/23-34t, z=23t\]

OpenStudy (stamp):

\[L(t)=<-25t+\frac{3}{23},-34t+\frac{5}{23},23t>\]

OpenStudy (stamp):

gg

OpenStudy (anonymous):

intersecting at right angles

OpenStudy (anonymous):

lag